OpenStudy (anonymous):

Help me please! Find an expression for the function whose graph is the given curve: the line segment joining the points (-3,-2) and (6,3)

OpenStudy (alekos):

well first you need to find the slope of the line segment. know how to do that?

OpenStudy (anonymous):

There's a formula for that, right?

OpenStudy (alekos):

sure is. slope = y-step/x-step = (y2-y1)/(x2-x1)

OpenStudy (anonymous):

Oh okay, so the slope would be 5/9, right?

OpenStudy (alekos):

you got it! now we need to find the y-intercept "b" using the slope and one of the points y = mx + b and m=slope

OpenStudy (anonymous):

So y= 5/9x + b, and then how do I find b?

OpenStudy (anonymous):

@alekos